you are preparing a solution and the recipe calls for 500mL of 2.00% glucose. what volume of 2.00M glucose stock would you need to make 5.00 mL of 2.00% glucose
Answer #1

We need 500 ml of 2% glucose = 2 g per 100 ml of water = 10 g glucose in 500 ml

Moles of glucose = mass of glucose / molar mass = 10 / 180 = 0.0556 mol

To find the volume of 2 M glucose nneded for this: Moles = molarity * Volume

0.0556 = 2 * V

Solving, V = 0.0278 Liters = 27.8 ml

Therefore we take 27.8 ml of 2 M solution and dilute it to 500 ml